# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

### 30 Mar 2026 v1

@@ Abstract
Orthogonalized-update optimizers such as Muon improve training of matrix-valued parameters, but existing extensions mostly act either after orthogonalization by rescaling updates or before it with heavier whitening-based preconditioners. We introduce MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon in three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). These variants rebalance the momentum matrix before finite-step Newton–Schulz using row/column squared-norm statistics and only $\mathcal{O}(m+n)$ auxiliary state. We show that finite-step orthogonalization is governed by input spectral properties, especially stable rank and condition number, and that row/column normalization is a zeroth-order whitening surrogate that removes marginal scale mismatch. For the hidden matrix weights targeted by MuonEq, the row-normalized variant R is the natural default and preserves the $\widetilde{\mathcal{O}}(T^{-1/4})$ stationarity guarantee of Muon-type methods. In LLaMA2 pretraining on C4, the default R variant consistently outperforms Muon on 130M and 350M models, yielding faster convergence and lower validation perplexity.
@@ 1 Introduction
As the computational and data costs of large language model pretraining grow, optimization increasingly determines sample efficiency, throughput, and scalability. Beyond coordinate-wise adaptive methods such as AdamW (Loshchilov and Hutter, 2019), Adan (Xie et al., 2024), and MARS (Liu et al., 2025b), recent work has shifted toward a geometric view of matrix-valued updates. Muon (Jordan et al., ), for instance, uses a Newton–Schulz approximation to the polar factor of the momentum matrix to produce spectrally aligned updates (Bernstein and Newhouse, 2024; Chen et al., 2025). It has shown strong empirical scalability in LLM pretraining, often improving the trade-off among data efficiency, compute, and wall-clock time at fixed target loss (Liu et al., 2025a; Shah et al., 2025). Recent theory further characterizes Muon as spectral-norm steepest descent and establishes nonconvex convergence guarantees for both exact and approximate orthogonalization (Chang et al., 2025; Sato et al., 2025; Li and Hong, 2025; Shen et al., 2025; Sfyraki and Wang, 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025).
Follow-up work around Muon mainly proceeds along two directions. (i) Post-orthogonalization rescaling augments an already constructed orthogonal update with additional magnitude correction. Muon+ (Zhang et al., 2026b), NorMuon (Li et al., 2025), and AdaMuon (Si et al., 2025) respectively add row/column normalization, neuron-wise normalization, or element-wise variance adaptation after orthogonalization. These variants show that orthogonal updates remain compatible with later adaptive scaling, but they do not change the matrix that finite-step Newton–Schulz actually sees. (ii) Pre-orthogonalization geometric correction instead modifies the coordinate system before spectral optimization. SOAP (Vyas et al., 2024) connects Shampoo (Gupta et al., 2018) with Adam/Adafactor (Shazeer and Stern, 2018; Kingma and Ba, 2015) through a preconditioned basis, while FISMO (Xu et al., 2026a) and Mousse (Zhang et al., 2026c) further perform spectral optimization in whitened coordinates. These methods suggest that improving geometry before orthogonalization is the more direct lever, but doing so usually requires maintaining left and right preconditioners together with extra matrix operations such as eigendecomposition, inverse power iteration, and whitening/dewhitening maps.
@@
Based on this observation, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon. MuonEq has three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). All three act directly on the momentum matrix before Newton–Schulz using only row/column squared-norm statistics and $\mathcal{O}(m+n)$ auxiliary state. Detailed justification is deferred to Section 3: Theorem 3.1 quantifies the spectral sensitivity of finite-step orthogonalization, Proposition 3.2 shows that row/column normalization is a zeroth-order surrogate for whitening, and Proposition 3.4 together with Theorem 3.5 clarify the trade-off between stronger two-sided equilibration and cleaner one-sided geometry. For the hidden Transformer weight matrices targeted by MuonEq, recent architecture-aware analyses by (Xu et al., 2026b) identify row-sided geometry as the natural one-sided choice, whereas column-sided geometry is more closely associated with embedding matrices; accordingly, we adopt R as the default and treat RC and C as companion forms for analysis and ablation. Compared with Muon+ (Zhang et al., 2026b), which rescales the update after orthogonalization, MuonEq modifies the matrix fed into orthogonalization. Unlike whitening-based methods such as FISMO and Mousse (Xu et al., 2026a; Zhang et al., 2026c), it avoids Kronecker factors, matrix-valued second-order statistics, eigendecomposition, and momentum transport, and requires only $\mathcal{O}(m+n)$ auxiliary state.
Our contributions are threefold.
  1. 1.
    First, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: RC, R, and C; these variants rebalance the matrix before Newton–Schulz using only row/column squared-norm statistics, with R as the default for hidden matrix weights.
  2. 2.
    Second, we show that finite-step Newton–Schulz is governed by the input spectrum, especially stable rank and condition number; that row/column normalization acts as a zeroth-order surrogate for whitening; and that RC and R admit a clear trade-off between stronger spectral correction and cleaner one-sided convergence structure, with the default R variant satisfying a $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee under standard assumptions.
  3. 3.
    Third, empirically, the default R variant consistently outperforms Muon across model sizes and token budgets, while ablations over RC, R, and C show that R is the strongest default form for the hidden-weight setting targeted by MuonEq.
@@ 2 Method
Notation and setup. Scalars are denoted by non-bold letters, vectors by bold lowercase letters, and matrices by bold uppercase letters. On $\mathbb{R}^{d}$, we use the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\operatorname{tr}(\mathbf{A}^{\top}\mathbf{B})$ and $\|\mathbf{A}\|_{F}$ denote the Frobenius inner product and norm, and $\|\mathbf{A}\|_{*}$ denotes the nuclear norm. We write $m]:=\{1,2,\ldots,m\}$ and let $\mathbb{N}$ be the set of non-negative integers. The model parameter is a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, and, unless otherwise stated, we assume $m\geq n$. For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with rank $r\leq n$, we write its compact singular value decomposition as $\mathbf{A}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$, and define $\operatorname{Orth}(\mathbf{A}):=\mathbf{U}\mathbf{V}^{\top}$. In finite-data form, the training objective can be written as $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $f_{i}(\mathbf{X})$ is the loss associated with the $i$-th sample $\mathbf{z}_{i}$. Equivalently, we use the stochastic formulation
@@
where $f$ may be nonconvex, $\xi$ is independent of $\mathbf{X}$, and $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$. For any matrix $\mathbf{A}$ with singular values $\{\sigma_{i}(\mathbf{A})\}_{i=1}^{r}$, we further use $\operatorname{sr}(\mathbf{A}):=\frac{\|\mathbf{A}\|_{F}^{2}}{\|\mathbf{A}\|_{2}^{2}},\kappa_{2}(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{r}(\mathbf{A})},p_{i}:=\frac{\sigma_{i}(\mathbf{A})^{2}}{\|\mathbf{A}\|_{F}^{2}},H(p):=-\sum_{i=1}^{r}p_{i}\log p_{i}.$
@@
$\mathbf{U}_{t}=\operatorname{Orth}(\tilde{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (2)
where $\tilde{\mathbf{M}}_{t}$ is either the standard momentum or its Nesterov variant (Jordan et al., ; Liu et al., 2025a; Chang et al., 2025). In practice, $\operatorname{Orth}(\tilde{\mathbf{M}}_{t})$ is implemented by a small fixed number of Newton–Schulz iterations rather than exact polar decomposition (Jordan et al., ). The design question is therefore not only what direction is desirable in the exact limit, but also what matrix should be fed into finite-step Newton–Schulz.
@@
$\hat{\mathbf{M}}_{t}=\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s),\qquad\mathbf{U}_{t}=\operatorname{Orth}(\hat{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (3)
@@
$\mathbf{D}_{r,t}=\operatorname{diag}\bigl(\operatorname{rowsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr),\qquad\mathbf{D}_{c,t}=\operatorname{diag}\bigl(\operatorname{colsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr).$
@@
$\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s)=\begin{cases}\mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{RC},\\ \mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t},&s=\mathrm{R},\\ \tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{C}.\end{cases}$ (4)
Since these maps use only row and column squared norms, the extra state is $\mathcal{O}(m+n)$, in the factorized spirit of Adafactor (Shazeer and Stern, 2018). Unlike methods that rely on matrix-valued or adaptive preconditioners (Gupta et al., 2018; Shazeer and Stern, 2018; Kingma and Ba, 2015; Xu et al., 2026a), MuonEq changes only the matrix fed to Newton–Schulz. Relative to Muon (Jordan et al., ), the orthogonalization routine itself is unchanged; relative to Muon+, NorMuon, and AdaMuon (Zhang et al., 2026b; Li et al., 2025; Si et al., 2025), MuonEq rebalances the input rather than rescales the output.
@@
$\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}\leq$ $\underbrace{\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\right\|_{F}}_{\text{finite-step approximation error}}$ (5)
$+\underbrace{\left\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}}_{\text{preconditioning bias}},$
@@
Unless otherwise stated, we use the row-normalized variant $s=\mathrm{R}$. The RC and C modes are kept as equally lightweight companion forms for analysis and ablation.
@@ 3 Analysis
The analysis follows the logic of MuonEq: Section 3.1 establishes the spectrum sensitivity of finite-step Newton–Schulz; Section 3.2 interprets row/column normalization as a zeroth-order surrogate for whitening; and Section 3.3 motivates the R rather than RC, with row normalization admitting a clean stochastic nonconvex guarantee. Together, Theorem 3.1, Proposition 3.2, Proposition 3.4, and Theorem 3.5 justify the MuonEq family and the default choice of R.

### 10 May 2026 v2

@@ Abstract
Orthogonalized-update optimizers such as Muon improve training of matrix-valued parameters, but existing extensions typically either rescale updates after orthogonalization or use heavier whitening-based preconditioners before it. We introduce MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). By rebalancing the momentum matrix before finite-step Newton–Schulz orthogonalization, MuonEq improves the geometry seen by orthogonalization. We show that finite-step orthogonalization is governed by the input spectrum, especially stable rank and condition number, and that row/column normalization acts as a zeroth-order surrogate for whitening. For hidden matrix weights, R is the default variant. Theoretically, MuonEq (R) retains the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ Muon-type nonconvex stationarity guarantee with decoupled weight decay and a horizon-free diminishing learning-rate schedule, and extends it to finite-step NS5 up to an explicit inexactness constant. In LLaMA2 pretraining on C4, MuonEq (R) consistently outperforms Muon on 130M, 350M, and 1B models, with faster convergence and lower validation perplexity. The code is available at the MuonEq codebase.
@@ 1 Introduction
As the computational and data costs of large language model pretraining grow, optimization increasingly determines sample efficiency, throughput, and scalability. Beyond coordinate-wise adaptive methods such as AdamW (Loshchilov and Hutter, 2019), Adan (Xie et al., 2024), and MARS (Yuan et al., 2024), recent work has shifted toward a geometric view of matrix-valued updates. Muon (Jordan et al., ), for instance, uses a Newton–Schulz approximation to the polar factor of the momentum matrix to produce spectrally aligned updates (Bernstein and Newhouse, 2024; Chen et al., 2025). It has shown strong empirical scalability in LLM pretraining, often improving the trade-off among data efficiency, compute, and wall-clock time at fixed target loss (Liu et al., 2025a; Shah et al., 2025; DeepSeek-AI, 2026). Recent theory further characterizes Muon as spectral-norm steepest descent and establishes nonconvex convergence guarantees for both exact and approximate orthogonalization (Chang et al., 2025; Sato et al., 2025; Li and Hong, 2025; Shen et al., 2025; Sfyraki and Wang, 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025).
Follow-up work around Muon mainly proceeds along two directions. (i) Post-orthogonalization rescaling augments an already constructed orthogonal update with additional magnitude correction. Muon+ (Zhang et al., 2026b), NorMuon (Li et al., 2025), and AdaMuon (Si et al., 2025) respectively add row/column normalization, neuron-wise normalization, or element-wise variance adaptation after orthogonalization. These variants show that orthogonal updates remain compatible with later adaptive scaling, but they do not change the matrix that finite-step Newton–Schulz actually sees. (ii) Pre-orthogonalization geometric correction instead modifies the coordinate system before spectral optimization. SOAP (Vyas et al., 2024) connects Shampoo (Gupta et al., 2018) with Adam/Adafactor (Shazeer and Stern, 2018; Kingma and Ba, 2015) through a preconditioned basis, while FISMO (Xu et al., 2026a) and Mousse (Zhang et al., 2026c) further perform spectral optimization in whitened coordinates. These methods motivate studying geometry correction before orthogonalization as a promising lever, but existing approaches typically incur substantially higher state and computational overhead.
@@
Motivated by this observation, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon, instantiated as two-sided row/column normalization (RC), row normalization (R), and column normalization (C). All variants rescale the momentum matrix before Newton–Schulz, using row/column squared-norm reductions from the current momentum. Section 3 provides the theory: Theorem 3.1 quantifies the spectral sensitivity of finite-step orthogonalization; Proposition 3.2 shows that row/column normalization is a zeroth-order whitening surrogate; and Proposition 3.4, Theorem 3.5, and Corollary 3.6 formalize the RC/R trade-off. In particular, the default R variant satisfies the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex stationarity guarantee while explicitly accounting for decoupled weight decay, horizon-free learning-rate schedules, and finite-step NS5. For the hidden Transformer weight matrices targeted by MuonEq, recent architecture-aware analyses (Xu et al., 2026b) identify row-sided geometry as the natural one-sided choice, while column-sided geometry is more aligned with embeddings; hence we use R by default and keep RC/C for analysis and ablation. Compared with Muon+ (Zhang et al., 2026b), which rescales updates after orthogonalization, MuonEq changes the input to orthogonalization. Unlike whitening-based methods such as FISMO and Mousse (Xu et al., 2026a; Zhang et al., 2026c), MuonEq avoids Kronecker factors, matrix-valued second-order statistics, eigendecomposition, and momentum transport; its diagonal rescaling is computed on the fly from row/column squared norms, with no persistent optimizer state.
Our contributions are threefold. (i) First, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: RC, R, and C; these variants rebalance the current momentum matrix before Newton–Schulz using row/column squared norms computed on the fly, with R as the default for hidden matrix weights. (ii) Second, we analyze how pre-orthogonalization geometry affects finite-step Newton–Schulz: Theorem 3.1 links its behavior to stable rank and condition number, Proposition 3.2 identifies row/column normalization as a zeroth-order whitening surrogate, and Proposition 3.4 together with Theorem 3.5 and Corollary 3.6 formalize the trade-off between stronger two-sided spectral correction and cleaner one-sided geometry. In particular, MuonEq (R) satisfies the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex stationarity guarantee while accounting for decoupled weight decay, horizon-free learning-rate schedules, and finite-step NS5. (iii) Third, empirically, the default R variant consistently outperforms Muon across model sizes and token budgets, while ablations over RC, R, and C show that R is the default form for the hidden-weight setting targeted by MuonEq.
@@ 2 Method
Notation and setup. Scalars are denoted by non-bold letters, vectors by bold lowercase letters, and matrices by bold uppercase letters. On $\mathbb{R}^{d}$, we use the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\operatorname{tr}(\mathbf{A}^{\top}\mathbf{B})$ and $\|\mathbf{A}\|_{F}$ denote the Frobenius inner product and norm, and $\|\mathbf{A}\|_{*}$ denotes the nuclear norm. We write $m]:=\{1,2,\ldots,m\}$ and let $\mathbb{N}$ be the set of non-negative integers. The model parameter is a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, and, unless otherwise stated, we assume $m\geq n$. For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with rank $r\leq n$, we write its compact singular value decomposition as $\mathbf{A}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$, and define the polar factor, which we also call the orthogonalized form, as $\mathrm{Orth}(\mathbf{A}):=\mathbf{U}\mathbf{V}^{\top}$. Here, orthogonalization and polar step mean applying or approximating this map, whereas whitening denotes full Gram inverse-square-root preconditioning. The row and column rescalings introduced below are diagonal equilibration operations. In finite-data form, the training objective can be written as $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $f_{i}(\mathbf{X})$ is the loss associated with the $i$-th sample $\mathbf{z}_{i}$. Equivalently, we use the stochastic formulation
@@
where $f$ may be nonconvex, $\xi$ is independent of $\mathbf{X}$, and $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$. For any matrix $\mathbf{A}$ with singular values $\{\sigma_{i}(\mathbf{A})\}_{i=1}^{r}$, we further use $\operatorname{sr}(\mathbf{A}):=\frac{\|\mathbf{A}\|_{F}^{2}}{\|\mathbf{A}\|_{2}^{2}},\kappa(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{r}(\mathbf{A})},\kappa_{i}(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{i}(\mathbf{A})},p_{i}:=\frac{\sigma_{i}(\mathbf{A})^{2}}{\|\mathbf{A}\|_{F}^{2}},H(p):=-\sum_{i=1}^{r}p_{i}\log p_{i}.$
@@
$\mathbf{O}_{t}=\operatorname{Orth}(\widetilde{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$ (2)
where $\widetilde{\mathbf{M}}_{t}$ is the standard momentum or its Nesterov variant (Jordan et al., ; Liu et al., 2025a; Chang et al., 2025); following LLM Muon practice (Liu et al., 2025a), we set $a=0.2\sqrt{\max\{m,n\}}$. In practice, $\operatorname{Orth}(\widetilde{\mathbf{M}}_{t})$ is implemented by a fixed number of Newton–Schulz iterations rather than exact polar decomposition (Jordan et al., ).
@@
$\hat{\mathbf{M}}_{t}=\textbf{DiagPre}(\widetilde{\mathbf{M}}_{t},s,\varepsilon),\qquad\mathbf{O}_{t}=\operatorname{Orth}(\hat{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$ (3)
@@
$\mathbf{D}_{r,t}=\operatorname{diag}\bigl(\operatorname{rowsum}(\widetilde{\mathbf{M}}_{t}\odot\widetilde{\mathbf{M}}_{t})+\varepsilon\bigr),\qquad\mathbf{D}_{c,t}=\operatorname{diag}\bigl(\operatorname{colsum}(\widetilde{\mathbf{M}}_{t}\odot\widetilde{\mathbf{M}}_{t})+\varepsilon\bigr).$
@@
$\textbf{DiagPre}(\widetilde{\mathbf{M}}_{t},s,\varepsilon)=\begin{cases}\mathbf{D}_{r,t}^{-1/2}\widetilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{RC},\\ \mathbf{D}_{r,t}^{-1/2}\widetilde{\mathbf{M}}_{t},&s=\mathrm{R},\\ \widetilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{C}.\end{cases}$ (4)
Unlike methods that rely on matrix-valued or adaptive preconditioners (Gupta et al., 2018; Shazeer and Stern, 2018; Kingma and Ba, 2015; Xu et al., 2026a), MuonEq changes only the transient input fed to Newton–Schulz. Relative to Muon (Jordan et al., ), the orthogonalization routine and persistent optimizer state are unchanged; relative to Muon+, NorMuon, and AdaMuon (Zhang et al., 2026b; Li et al., 2025; Si et al., 2025), MuonEq rebalances the input rather than rescales the output.
@@
$\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}\leq\underbrace{\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\right\|_{F}}_{\text{finite-step approximation error}}+\underbrace{\left\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}}_{\text{preconditioning bias}},$ (5)
@@
Unless otherwise stated, we use $s=\mathrm{R}$; $\mathrm{RC}$ and $\mathrm{C}$ are retained for analysis and ablation. For varying $\beta_{t}$, the Nesterov line follows the EMA-buffer convention; for $\beta_{t}\equiv\beta$, it reduces to the Nesterov form Yuan et al. (2024); Chang et al. (2025).
@@ 3 Analysis
The analysis follows the logic of MuonEq: Section 3.1 establishes the spectrum sensitivity of finite-step Newton–Schulz; Section 3.2 interprets row/column normalization as a zeroth-order surrogate for whitening; and Section 3.3 motivates the R rather than RC, with row normalization admitting a clean stochastic nonconvex guarantee. Together, these results motivate the MuonEq family and clarify the trade-off between stronger spectral correction and preprocessing bias. Within the hidden-weight setting studied here, they support using R as our default working variant.